OpenStudy (anonymous):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)? x − y = 1 2x + 2y = 4 x + y = 1 x + 2y = 2 Please just answer (:

OpenStudy (anonymous):

Yes each equation: let x=0, and y=1

OpenStudy (anonymous):

ex: x-y=1 0-1=1? -1=1? no

OpenStudy (anonymous):

test each to find the one which is true

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

no, n=n to be true

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

x+y=1 0+1=1 1=1 yes

OpenStudy (anonymous):

thank you!